Write the transformed function f(x) = 8^x is horizontally compressed by a factor of 2.
How would i slove this?
In general functions are compressed horizontally by a factor of two when you replace the 'x' in the function with '2x'.
The solution for this particular problem would be:
f(x) = 8 ^(2x)
This can be simplified...
f(x) = 8 ^ (2 * x)
f(x) = (8 ^ 2) ^ x
f(x) = 64 ^ x


Answer by jsmallt9(3759)

In general functions are compressed horizontally by a factor of two when you replace the 'x' in the function with '2x'.
The solution for this particular problem would be:
f(x) = 8 ^(2x)
This can be simplified...
f(x) = 8 ^ (2 * x)
f(x) = (8 ^ 2) ^ x
f(x) = 64 ^ x